OpenStudy (anonymous):

what is the solutions for y ≥ x4 – 1

OpenStudy (anonymous):

\[y \ge (x^2-1)(x^2+1)\] \[y \ge (x-1)(x+1)(x^2+1)\] Sketch the diagram. 

OpenStudy (anonymous):

Sketch it and post it here to let me see.

OpenStudy (anonymous):

is it -1<x<1

OpenStudy (anonymous):

Correct. Awesome job. Well done.

OpenStudy (anonymous):

Wait it's greater than or equal sign

OpenStudy (anonymous):

@514sophie

OpenStudy (anonymous):

\[-1 \le x \le 1\]